One number is 4 less than three times a number. If the product of these two numbers is 84, find the numbers.

Answer by checkley79(3341)

X=3Y-4
XY=84
(3Y-4)Y=84
3Y^2-4Y-84=0
(3Y-14)(Y-6)=0
Y-6=0
Y=6 ANS.
X=3*6-4
X=18-4=14 ANS.
PROOF:
6*14=84
84=84